# Robust Decision Transformer: Tackling Data Corruption in Offline RL via Sequence Modeling

### v1

@@ D.6 Detailed Results for the Main Experiments
Table 5: Detailed performance under random data corruption.
Environment Attack Element BC RBC DeFog CQL UWMSG RIQL DT RDT (ours) halfcheetah state 33.7$\pm$0.6 33.0$\pm$2.7 15.8$\pm$3.8 16.7$\pm$1.2 1.1$\pm$0.8 19.9$\pm$2.1 29.7$\pm$2.2 28.1$\pm$0.8 action 34.9$\pm$2.9 35.6$\pm$0.5 34.2$\pm$2.4 44.5$\pm$0.3 49.9$\pm$0.7 41.3$\pm$1.4 36.6$\pm$0.9 35.7$\pm$1.0 reward 36.0$\pm$1.2 34.8$\pm$2.5 33.5$\pm$3.7 44.1$\pm$0.3 37.2$\pm$5.8 42.7$\pm$0.9 38.9$\pm$0.6 37.7$\pm$1.1 hopper state 25.4$\pm$16.0 33.1$\pm$8.7 24.0$\pm$9.0 21.8$\pm$5.9 16.4$\pm$5.0 34.0$\pm$13. 53.7$\pm$11.6 64.6$\pm$6.4 action 25.3$\pm$11.8 33.5$\pm$12.2 52.9$\pm$17.8 55.4$\pm$7.9 49.8$\pm$12.2 59.6$\pm$12.7 59.4$\pm$1.7 66.0$\pm$5.6 reward 28.2$\pm$5.6 33.9$\pm$13.2 45.1$\pm$13.1 57.2$\pm$14.4 65.4$\pm$26.0 59.6$\pm$4.8 57.7$\pm$11.6 63.8$\pm$11.3 walker2d state 13.1$\pm$8.0 28.5$\pm$7.6 18.5$\pm$3.5 27.6$\pm$10.2 -0.5$\pm$0.3 14.2$\pm$1.2 53.2$\pm$10.2 54.5$\pm$2.3 action 19.6$\pm$9.2 19.6$\pm$9.2 61.9$\pm$2.8 70.8$\pm$5.3 57.1$\pm$10.4 72.6$\pm$17.3 61.2$\pm$6.7 60.6$\pm$2.9 reward 20.1$\pm$4.8 18.6$\pm$4.7 50.2$\pm$14.8 74.6$\pm$3.0 51.3$\pm$10.6 79.0$\pm$5.8 61.4$\pm$6.5 62.9$\pm$5.6 halfcheetah (10%) state 5.2$\pm$1.6 1.5$\pm$0.2 3.9$\pm$1.0 11.3$\pm$3.4 5.4$\pm$1.2 4.4$\pm$0.9 7.0$\pm$0.5 8.0$\pm$0.5 action 1.8$\pm$0.2 0.6$\pm$0.1 8.7$\pm$1.6 28.4$\pm$3.2 11.7$\pm$3.4 2.3$\pm$0.6 4.3$\pm$0.3 13.8$\pm$2.3 reward 4.4$\pm$0.8 1.9$\pm$0.2 12.4$\pm$1.7 34.2$\pm$1.0 2.2$\pm$0.7 6.1$\pm$1.2 12.7$\pm$1.0 16.3$\pm$2.4 hopper (10%) state 15.0$\pm$0.6 14.4$\pm$2.0 15.6$\pm$6.5 2.0$\pm$0.4 15.6$\pm$8.0 15.5$\pm$5.4 37.3$\pm$1.1 40.0$\pm$3.2 action 19.2$\pm$3.6 17.7$\pm$1.2 23.7$\pm$4.8 1.8$\pm$0.0 40.7$\pm$5.9 26.3$\pm$2.8 32.1$\pm$1.9 37.9$\pm$7.9 reward 15.1$\pm$0.7 21.1$\pm$0.7 19.2$\pm$10.2 2.6$\pm$0.9 18.2$\pm$2.8 40.2$\pm$2.6 39.9$\pm$2.8 44.3$\pm$4.7 walker2d (10%) state 8.2$\pm$1.8 8.1$\pm$0.9 10.6$\pm$2.2 0.3$\pm$0.7 -0.5$\pm$0.4 9.2$\pm$4.4 19.9$\pm$3.4 21.3$\pm$1.9 action 6.1$\pm$1.0 4.0$\pm$1.6 13.7$\pm$4.7 0.7$\pm$0.9 5.6$\pm$1.2 9.8$\pm$1.9 21.9$\pm$2.7 25.9$\pm$2.0 reward 10.4$\pm$1.6 9.9$\pm$1.5 17.0$\pm$2.5 1.2$\pm$1.6 7.9$\pm$0.8 9.6$\pm$2.5 29.8$\pm$0.4 34.1$\pm$5.3 kitchen-complete state 47.2$\pm$3.3 11.6$\pm$10.9 40.1$\pm$8.5 15.1$\pm$10.4 0.4$\pm$0.6 10.2$\pm$6.2 43.8$\pm$3.6 52.0$\pm$2.8 action 16.1$\pm$9.5 16.6$\pm$11.3 16.5$\pm$15.4 10.9$\pm$6.7 0.0$\pm$0.0 0.2$\pm$0.2 20.9$\pm$2.6 43.8$\pm$2.4 reward 38.8$\pm$18.6 56.9$\pm$8.9 41.8$\pm$10.6 0.0$\pm$0.0 0.0$\pm$0.0 25.0$\pm$15.9 65.8$\pm$4.6 71.2$\pm$2.9 kitchen-partial state 34.1$\pm$1.6 32.4$\pm$3.0 11.1$\pm$10.5 0.1$\pm$0.2 0.0$\pm$0.0 26.0$\pm$4.7 27.4$\pm$3.5 40.4$\pm$9.6 action 43.8$\pm$1.6 44.5$\pm$4.4 12.8$\pm$8.6 0.0$\pm$0.0 0.0$\pm$0.0 39.2$\pm$7.0 32.4$\pm$2.1 32.5$\pm$2.6 reward 42.8$\pm$7.8 39.4$\pm$7.2 8.2$\pm$10.0 0.4$\pm$0.6 0.0$\pm$0.0 53.6$\pm$12.8 39.2$\pm$3.0 42.4$\pm$6.4 kitchen-mixed state 38.2$\pm$12.0 45.5$\pm$4.3 11.1$\pm$5.0 0.0$\pm$0.0 0.0$\pm$0.0 31.2$\pm$4.9 22.5$\pm$11.4 36.9$\pm$7.3 action 53.9$\pm$1.8 50.9$\pm$3.4 11.8$\pm$6.5 0.0$\pm$0.0 0.0$\pm$0.0 52.4$\pm$1.9 33.1$\pm$2.7 44.9$\pm$4.2 reward 54.4$\pm$2.6 54.6$\pm$2.3 9.1$\pm$8. 0.0$\pm$0.0 0.0$\pm$0.0 50.4$\pm$2.9 42.8$\pm$2.2 44.1$\pm$5.6 door (1%) state 74.8$\pm$18.2 96.5$\pm$7.1 100.7$\pm$3.4 -0.3$\pm$0.0 -0.2$\pm$0.2 31.3$\pm$19.6 99.7$\pm$3.7 103.7$\pm$0.7 action 5.3$\pm$3.4 79.0$\pm$9.0 104.0$\pm$0.9 -0.3$\pm$0.0 -0.2$\pm$0.1 11.5$\pm$11.9 85.9$\pm$2.9 104.8$\pm$0.2 reward 22.5$\pm$13.6 96.2$\pm$3.0 102.5$\pm$0.9 -0.3$\pm$0.0 -0.2$\pm$0.1 17.5$\pm$13.4 94.4$\pm$5.9 104.3$\pm$0.5 hammer (1%) state 87.9$\pm$34.9 95.8$\pm$21.3 96.6$\pm$16.8 0.3$\pm$0.0 0.0$\pm$0.2 35.2$\pm$8.4 99.3$\pm$8.7 120.8$\pm$3.0 action 92.8$\pm$17.9 97.3$\pm$11.3 89.9$\pm$13.3 0.3$\pm$0.0 0.3$\pm$0.1 25.6$\pm$13.6 90.3$\pm$8.2 111.1$\pm$5.3 reward 72.5$\pm$23.7 112.6$\pm$11.7 99.5$\pm$19.2 0.3$\pm$0.0 0.3$\pm$0.1 47.0$\pm$5.0 94.9$\pm$16.8 117.3$\pm$8.3 relocate (1%) state 0.9$\pm$0.6 30.3$\pm$17.2 21.5$\pm$1.8 -0.2$\pm$0.1 -0.3$\pm$0.0 2.6$\pm$3.5 61.7$\pm$5.1 64.7$\pm$4.0 action 9.6$\pm$10.2 39.2$\pm$10.4 57.8$\pm$4.9 -0.3$\pm$0.1 -0.3$\pm$0.0 0.7$\pm$0.6 54.2$\pm$4.1 81.5$\pm$4.7 reward 9.1$\pm$5.6 57.1$\pm$16.1 64.5$\pm$4.6 -0.3$\pm$0.1 -0.3$\pm$0.0 3.5$\pm$2.7 64.8$\pm$4.1 81.1$\pm$5.4
Table 6: Detailed performance under adversarial data corruption.
Environment Attack Element BC RBC DeFog CQL UWMSG RIQL DT RDT (ours) halfcheetah (10%) state 1.9$\pm$0.1 1.8$\pm$0.3 3.4$\pm$0.5 9.9$\pm$0.6 4.9$\pm$1.5 3.6$\pm$0.5 7.4$\pm$0.6 7.5$\pm$0.4 action -0.2$\pm$0.2 -0.4$\pm$0.2 7.8$\pm$1.4 32.3$\pm$1.1 6.0$\pm$0.4 0.8$\pm$0.4 4.1$\pm$0.5 12.1$\pm$1.1 reward 4.4$\pm$0.8 1.8$\pm$0.4 13.0$\pm$1.3 32.5$\pm$2.5 1.9$\pm$0.2 10.9$\pm$1.6 13.3$\pm$1.6 16.9$\pm$3.3 hopper (10%) state 23.2$\pm$3.3 22.1$\pm$3.2 12.5$\pm$3.2 2.5$\pm$0.1 15.1$\pm$4.1 19.1$\pm$6.9 38.6$\pm$4.7 39.3$\pm$5.1 action 14.9$\pm$1.6 17.3$\pm$3.3 14.5$\pm$10.3 1.5$\pm$0.5 42.6$\pm$3.6 17.1$\pm$1.4 26.9$\pm$5.7 27.4$\pm$2.0 reward 15.1$\pm$0.7 22.2$\pm$2.8 16.4$\pm$9.6 1.8$\pm$0.0 28.6$\pm$9.9 34.5$\pm$5.1 40.5$\pm$5.5 40.8$\pm$4.8 walker2d (10%) state 7.2$\pm$0.6 8.6$\pm$0.6 10.3$\pm$5.0 1.2$\pm$1.3 3.3$\pm$2.3 10.4$\pm$1. 22.3$\pm$2.4 21.1$\pm$2.6 action 2.0$\pm$0.7 2.2$\pm$1.0 8.8$\pm$3.2 0.5$\pm$0.8 5.9$\pm$0.8 3.3$\pm$1.2 7.7$\pm$0.8 13.0$\pm$1.4 reward 10.4$\pm$1.6 8.4$\pm$1.5 12.2$\pm$6.5 0.9$\pm$1.0 9.7$\pm$3.1 9.3$\pm$1.2 29.2$\pm$3.6 29.8$\pm$5.3 kitchen-complete state 15.8$\pm$12.8 25.4$\pm$19.0 40.1$\pm$8.5 5.2$\pm$9.1 0.0$\pm$0.0 5.9$\pm$5.0 48.4$\pm$6.7 58.4$\pm$3.7 action 11.5$\pm$4.4 10.0$\pm$6.6 16.5$\pm$15.4 0.1$\pm$0.2 0.0$\pm$0.0 1.1$\pm$1.1 21.1$\pm$2.3 24.8$\pm$3.1 reward 38.8$\pm$18.6 41.8$\pm$13.5 41.8$\pm$10.6 0.0$\pm$0.0 0.0$\pm$0.0 19.8$\pm$4.7 67.5$\pm$5.3 72.5$\pm$6.3 kitchen-partial state 38.4$\pm$4.3 36.8$\pm$4.3 11.1$\pm$10.5 0.0$\pm$0.0 0.0$\pm$0.0 34.9$\pm$2.8 32.6$\pm$6.1 36.5$\pm$8.8 action 2.6$\pm$1.6 6.1$\pm$6.1 12.8$\pm$8.6 1.9$\pm$3.2 9.4$\pm$10.4 7.5$\pm$2.7 2.9$\pm$2.2 27.9$\pm$2.7 reward 42.8$\pm$7.8 37.6$\pm$4.4 8.2$\pm$10.0 0.0$\pm$0.0 0.0$\pm$0.0 43.6$\pm$5.2 34.1$\pm$4.5 39.4$\pm$5.4 kitchen-mixed state 46.2$\pm$3.3 49.6$\pm$3.3 11.1$\pm$5.0 0.0$\pm$0.0 0.0$\pm$0.0 44.8$\pm$4.1 28.2$\pm$9.9 30.0$\pm$5.5 action 0.9$\pm$1.5 7.1$\pm$7.9 11.8$\pm$6.5 0.0$\pm$0.0 6.0$\pm$10.4 4.5$\pm$4.6 0.2$\pm$0.4 48.5$\pm$6.4 reward 54.4$\pm$2.6 50.8$\pm$3.3 9.1$\pm$8.1 0.0$\pm$0.0 0.0$\pm$0.0 52.9$\pm$3.1 47.9$\pm$2.8 50.1$\pm$5.0 door (1%) state 75.3$\pm$14.9 82.2$\pm$13.2 102.6$\pm$1.3 -0.3$\pm$0.0 -0.2$\pm$0.0 38.0$\pm$22.8 99.0$\pm$0.9 104.7$\pm$0.5 action 11.2$\pm$7.4 30.1$\pm$8.0 102.8$\pm$1.4 -0.3$\pm$0.1 -0.2$\pm$0.0 1.0$\pm$0.8 46.9$\pm$6.9 102.0$\pm$0.9 reward 22.5$\pm$13.6 90.9$\pm$3.8 103.0$\pm$0.8 -0.3$\pm$0.0 -0.2$\pm$0.0 28.5$\pm$8.1 95.7$\pm$3.3 104.1$\pm$0.6 hammer (1%) state 85.4$\pm$23.8 96.2$\pm$6.3 105.8$\pm$11.5 0.3$\pm$0.0 0.2$\pm$0.1 67.0$\pm$12.4 96.0$\pm$2.5 116.6$\pm$7.4 action 37.3$\pm$12.5 32.1$\pm$21.1 35.1$\pm$17.0 0.2$\pm$0.0 0.2$\pm$0.0 14.5$\pm$6.3 48.9$\pm$10.5 119.1$\pm$4.5 reward 72.5$\pm$23.7 107.7$\pm$16.0 80.5$\pm$12.5 0.3$\pm$0.0 0.1$\pm$0.1 47.8$\pm$19.7 97.1$\pm$7.0 118.3$\pm$7.1 relocate (1%) state 39.2$\pm$20.3 33.0$\pm$8.4 32.9$\pm$5.5 -0.3$\pm$0.0 -0.3$\pm$0.0 6.1$\pm$6.7 76.2$\pm$5.0 69.0$\pm$4.4 action 9.4$\pm$6.7 10.8$\pm$7.0 24.7$\pm$3.9 -0.2$\pm$0.0 -0.3$\pm$0.0 0.3$\pm$0.4 9.8$\pm$2.7 49.5$\pm$7.8 reward 9.1$\pm$5.6 53.4$\pm$20.0 65.2$\pm$8.3 -0.3$\pm$0.0 -0.3$\pm$0.0 3.0$\pm$3.1 70.8$\pm$4.5 51.8$\pm$7.0

### v3

@@ D.11 Performance under MuJoCo Dataset with Different Quality Levels
Table 8: Results on “medium-v2” dataset under random data corruption.
Attack Task DeFog RIQL DT RDT State halfcheetah (2%) 7.5$\pm$ 1.4 18.0$\pm$ 1.5 15.7$\pm$ 1.3 22.3$\pm$ 1.1 hopper (2%) 20.4$\pm$ 8.3 45.7$\pm$ 7.3 48.6$\pm$ 3.2 52.2$\pm$ 5.7 walker2d (2%) 14.2$\pm$ 1.6 25.4$\pm$ 5.0 20.5$\pm$ 6.0 28.0$\pm$ 8.0 Average 14.0 29.7 28.3 34.2 Action halfcheetah (2%) 24.3$\pm$ 1.4 13.9$\pm$ 1.9 16.2$\pm$ 0.8 32.8$\pm$ 1.2 hopper (2%) 36.7$\pm$ 9.6 51.9$\pm$ 3.9 37.5$\pm$ 6.0 45.3$\pm$ 3.2 walker2d (2%) 30.7$\pm$ 7.0 27.5$\pm$ 5.3 38.9$\pm$ 2.9 43.8$\pm$ 7.9 Average 30.5 31.1 30.9 40.6 Reward halfcheetah (2%) 27.4$\pm$ 2.9 36.1$\pm$ 1.3 33.3$\pm$ 1.0 37.5$\pm$ 2.0 hopper (2%) 32.5$\pm$ 17.5 55.5$\pm$ 6.8 51.6$\pm$ 2.7 53.9$\pm$ 4.6 walker2d (2%) 28.2$\pm$ 18.4 42.1$\pm$ 13.4 50.5$\pm$ 5.1 62.1$\pm$ 4.7 Average 29.3 44.6 45.1 51.2 Average over all tasks 24.6 35.1 34.8 42.0
Table 9: Results on “expert-v2” dataset under random data corruption.
Attack Task DeFog RIQL DT RDT State halfcheetah (2%) 3.8$\pm$ 2.3 0.5$\pm$ 1.2 2.9$\pm$ 0.8 4.4$\pm$ 0.2 hopper (2%) 15.1$\pm$ 4.2 32.0$\pm$ 4.4 38.5$\pm$ 7.4 48.6$\pm$ 7.0 walker2d (2%) 15.3$\pm$ 3.1 21.7$\pm$ 4.6 40.7$\pm$ 4.8 41.6$\pm$ 4.2 Average 11.4 18.1 27.4 31.5 Action halfcheetah (2%) 2.7$\pm$ 1.0 -1.3$\pm$ 1.5 3.2$\pm$ 2.4 5.4$\pm$ 0.7 hopper (2%) 18.0$\pm$ 4.7 32.4$\pm$ 8.7 21.2$\pm$ 3.3 36.0$\pm$ 4.8 walker2d (2%) 25.6$\pm$ 3.8 7.1$\pm$ 1.0 33.1$\pm$ 6.8 73.8$\pm$ 3.6 Average 15.5 12.7 19.2 38.4 Reward halfcheetah (2%) 2.5$\pm$ 0.8 2.7$\pm$ 0.9 4.9$\pm$ 0.7 17.2$\pm$ 5.5 hopper (2%) 21.2$\pm$ 6.3 89.5$\pm$ 4.6 49.5$\pm$ 3.3 67.3$\pm$ 6.6 walker2d (2%) 36.0$\pm$ 9.8 64.2$\pm$ 11.0 96.3$\pm$ 2.6 103.1$\pm$ 3.6 Average 19.9 52.1 50.2 62.5 Average over all tasks 15.6 27.6 32.3 44.1
D.12 Performance under NeoRL Dataset
To further demonstrate the robustness of the RDT, we evaluate it on the NeoRL benchmark (Qin et al., 2022). Specifically, we select the “citylearn-medium” and “finance-medium” datasets as the testbeds. The Finance dataset enables the construction of a trading simulator, while the CityLearn dataset reshapes the aggregation curve of electricity demand by controlling energy storage. We randomly sample $20\%$ and $10\%$ of the Finance and CityLearn datasets, respectively, resulting in both datasets having a similar size of approximately $2\times 10^{4}$ transitions. We consider RIQL, DeFog, and DT to be comparable baselines. The performance is measured using the response reward criterion, as detailed in Table 10. Notably, RDT achieves the best average performance compared to the other baselines, outperforming DT by 12.13% under Finance dataset. These results underscore RDT’s robustness and potential for real-world applications.
Table 10: Results on NeoRL dataset under random data corruption.
Task Attack DeFog RIQL DT RDT citylearn (20%) State 39138. 9 30235. 3 37878. 9 39124. 2 Action 38970. 6 29093. 2 35589. 0 39066. 4 Reward 38156. 4 29634. 8 37626. 9 38909. 8 Average 38755. 3 29654. 4 37031. 6 39033. 5 finance (10%) State 402. 2 339. 4 403. 4 411. 0 Action 386. 6 400. 6 352. 6 452. 1 Reward 330. 3 376. 3 329. 5 437. 6 Average 373. 0 372. 1 361. 8 433. 5
D.13 Additional Evaluation on Iterative Data Correction
We have demonstrated the effectiveness of iterative data correction by illustrating the MSE loss between corrected and original data in Figure 3(c). The MSE loss gradually decreases as training proceeds, indicating that RDT can accurately predict the correct data. To further investigate the effectiveness of iterative data correction, we record the precision of detection using the $z$-score. Specifically, DT w. IDC can detect $N$ data points as corrupted within a batch, and $M$ out of these $N$ data points are truly corrupted. We record the ratio of $M/N$ under action attack as shown in Figure 18. As observed, the detection precision increases with the training process and can reach 80% to 90% at the 50th epoch, demonstrating the effectiveness of IDC. It is worth noting that the precision is low at the start of training; therefore, we begin data correction at the 50th training epoch in our default implementation.
Figure 18: Accuracy of iterative data correction detection via the $z$-score.
D.14 Discussion on State Prediction and Correction
In the implementation of RDT, we predict both actions and rewards to tackle data corruption. We justify that reward prediction can provide benefits under state and action attacks (see Section D.2). Here, we explore the impact of state prediction. We create a variant, DT w. SP, based on the original DT, which additionally predicts states. We compare DT and DT w. SP under state and action attacks, as shown in Figure 19(a). As observed, state prediction does not consistently benefit all tasks. The performance drops on the Kitchen and Adroit tasks, perhaps due to the larger state dimensions. For example, the state dimension of Kitchen is 60, which is larger than its action dimension of 9.
Based on DT w. SP, we evaluate state correction via iterative data correction under different $\zeta$, as shown in Figure 19(b). The results indicate that state correction still causes performance drops. Indeed, state prediction involves learning the environment’s dynamics (i.e., transition probabilities), which is inherently challenging, as prediction errors in state prediction negatively impact policy learning. Additionally, in our experimental setting, data corruption exacerbates prediction errors, leading to further performance degradation. We plan to draw inspiration from model-based RL work to address this issue in future research.
(a) Ablation on state prediction.
(b) Ablation on state correction.
Figure 19: Ablation study on state prediction and correction. (a) Comparison of DT and DT with State Prediction (DT w. SP), showing that DT w. SP does not consistently bring improvements. (b) Evaluation of state correction under different $\zeta$ values. Results indicate that state prediction can cause performance drops.
D.15 Motivating Example under Different Data Corruption
To demonstrate the performance drop under varying dataset sizes from another perspective, we modify Figure 1 to create Figure 20. In Figure 20, we place the performances of different algorithms in separate boxes to better illustrate their sensitivity to different types of data corruption. As observed, UWMSG and RIQL are very sensitive to state attacks. This sensitivity arises because we do not separate state and next-state attacks but apply state attacks within a single trajectory, representing a more general setting. UWMSG and RIQL are algorithms based on temporal difference learning, which learns the $Q$-value dependent on the next state. Hence, they experience significant performance drops under state attacks in our settings. In contrast, DeFog and DT, which are based on sequence modeling, are comparatively robust to all types of data corruption. Furthermore, DT demonstrates better robustness than other methods, even with limited datasets. Therefore, we propose three robustness techniques to further improve robustness against data corruption.
Figure 20: An alternative version of the motivating example, where the performances of different algorithms are placed in separate boxes to better illustrate their sensitivity to various data corruption.
Appendix E Limitation and Discussion
One limitation of RDT is our decision to avoid state prediction and correction within the robust sequence modeling framework. The high dimensionality of states poses significant challenges to achieving accurate predictions. Additionally, data corruption can further complicate state prediction, leading to potential declines in performance, as shown in Section D.14. Despite this, we have validated the feasibility and effectiveness of robust sequence modeling in this paper, and we leave state prediction and correction as promising directions for future exploration.